# Estimating the growth rate of a birth and death process using data from a small sample

### v1

@@ 3.1 Simulation Results
3.1.1 Mean Squared Error and Mean Absolute Error
@@
We conducted a simulation study to evaluate our estimators $\hat{r}_{MSE}$, $\hat{r}_{Inv}$, and $\hat{r}_{Bias}$, using simulation scenarios similar to those in [7]. Specifically, we simulated coalescent point processes for $n=5\text{--}10,15,20,50$, with $T=40$ and $r=0.5,1$, repeating each scenario 10,000 times for each sample size and growth rate. The simulations were carried out using the R package cloneRate. Based on these simulations, we compared the mean squared error (MSE) and the mean absolute error (MAE) of the estimators $\hat{r}_{MSE},\hat{r}_{Bias},\hat{r}_{Inv},\hat{r}_{MCMC},\hat{r}_{Lengths}$, and $\hat{r}_{MLE}$. Here $\hat{r}_{MCMC}$ is the Markov chain Monte Carlo estimator based on the work of [14], which is implemented in cloneRate.
The mean squared errors for $r=0.5$ and $r=1$, with $T=40$, across the six different estimators are shown in Figures 3 and 3, respectively. These figures show that the MSE of the estimator $\hat{r}_{{MSE}}$ is the lowest across all sample sizes. The differences between estimators decrease as $n$ increases. For example, for $n=5$ and $r=0.5$, the MSE of $\hat{r}_{{MCMC}}$ is nearly four times that of $\hat{r}_{{MSE}}$, whereas for $n=15$ and $r=0.5$, the MSE of $\hat{r}_{{MCMC}}$ is only about twice that of $\hat{r}_{{MSE}}$. Moreover, not only does $\hat{r}_{{MSE}}$ outperform $\hat{r}_{{MCMC}},\hat{r}_{{MLE}}$ and $\hat{r}_{{Lengths}}$, but also $\hat{r}_{{Bias}}$ and $\hat{r}_{{Inv}}$ achieve lower MSEs than these previously existing methods.
Figure 2: Comparison of the MSE of the six estimators when $r=0.5$ and $T=40$.
Figure 3: Comparison of the MSE of the six estimators when $r=1$ and $T=40$.
@@
The mean absolute errors of the estimators are shown in Figure 5 and Figure 5 for $T=40$ and $r=0.5,r=1$. Here, $\hat{r}_{Bias}$ has the lowest mean absolute error across the six estimators. However, again, also $\hat{r}_{MSE}$ and $\hat{r}_{Inv}$ perform better than $\hat{r}_{MCMC},\hat{r}_{MSE}$ and $\hat{r}_{Lengths}.$ The differences among the estimators decrease for increasing $n.$ Moreover, $\hat{r}_{MSE}$, $\hat{r}_{Bias}$ and $\hat{r}_{Inv}$ outperform the previously existing approaches even for $n=50.$
@@
For $\hat{r}_{MCMC}$, we assumed a maximal growth rate of $r=4$, as this is the default value in the software package. If we applied the same assumption to the other estimators, it would further reduce their MSE and MAE. It is shown in Figure 14 in the appendix that the choice of the maximal growth rate substantially impacts the mean squared error of the estimator $\hat{r}_{MCMC}$.
3.1.2 Bias and densities of the estimators
The densities of the six estimators are shown in Figure 7 for $n=5$ and in Figure 7 for $n=20$. These plots show that even for $n=5$, the maximal points of the densities of $\hat{r}_{Lengths}$, $\hat{r}_{MLE}$, $\hat{r}_{MCMC}$, and $\hat{r}_{Inv}$ are similar to the true growth rate. According to the density plots, $\hat{r}_{Bias}$ and especially $\hat{r}_{MSE}$ typically underestimate the growth rate for small $n$. Moreover, the densities are all very skewed for $n=5$.
For $n=20,$ the differences among the estimators are much smaller than for $n=5.$ The skewness decreases with increasing sample size so that for $n=20$, the densities look similar to the density of a normal distribution. This aligns with results in [7].
The bias of the estimators is presented in Figures 9 and 9. According to the figure, the estimates $\hat{r}_{Lengths}$, $\hat{r}_{MLE}$, and $\hat{r}_{MCMC}$ all tend to overestimate the growth rate on average for small $n$. The estimate $\hat{r}_{Bias}$, which was designed to be exactly unbiased when $T=\infty$, remains nearly unbiased even when $T=40$. The estimate $\hat{r}_{Inv}$ has a slight upward bias, as would be expected from Jensen’s inequality because it is the reciprocal of an estimator which is unbiased for $1/r$ when $T=\infty$. The estimate $\hat{r}_{MSE}$ underestimates the growth rate on average when $n$ is small. It acts as a shrinkage estimator, minimizing the MSE by introducing a downward bias to reduce the variance and mitigate the impact of rare events in which the growth rate is vastly overestimated.
@@ 3.1.2 Bias and densities of the estimators
The density plots in Figure 7 suggest that the mean squared error is probably not the best criterion to use to evaluate these estimators, especially when $n$ is small. The mean squared error is substantially influenced by rare events when the methods estimate a huge growth rate because all of the coalescence times are very close to one another. Consequently, the estimator $\hat{r}_{MSE}$ that minimizes the mean squared error often underestimates $r$ by nearly a factor of $2$ when $n$ is small. The estimator $r_{Inv}$ appears to be a good compromise. It still achieves a lower mean absolute error than $\hat{r}_{MCMC}$ when $n\geq 6$, and the mode of the distribution is close to the true growth rate. It also has the benefit that $c_{Inv}(n)$ can be computed exactly.
@@ 3.1.4 Influence of the Approximation $T=\infty$ on the Quality of the Estimator
Now, we quantify the error introduced by this approximation. That is, we assess how a change in the constant influences the MSE and the bias and how the constant would differ if it were allowed to depend on $T$. Given $n$, $T$, and $r$, we can determine by simulation, for any value of the constant $c$, the mean squared error and the absolute bias of the resulting estimator of $r$. In Figures 12 and 12, we plot the MSE and the bias as a function of the constant.
Figure 11: Influence of the constant on the MSE for $T=40$, $n=10$, and $r=0.5$. The black, dotted line represents the true minimal value for the constant and the green dotted line represents $c_{MSE}(10).$ The purple dotted line represents $c_{Bias}(10).$
Figure 12: Influence of the constant on the MAE for $T=40$, $n=10$, and $r=0.5$. The black, dotted line represents the true minimal value for the constant and the green dotted line represents $c_{MSE}(10).$ The purple dotted line represents $c_{Bias}(10).$
In the figures, the green and purple lines do not exactly align with the minimum points of the MSE and the absolute bias, respectively. This discrepancy arises because the values for these lines were calculated under the assumption that $T=\infty$, whereas the data for the blue curves were simulated with $T=40$. However, we observe that the green dotted line in Figure 12 lies close to the minimum point of the curve, which means the difference between the constant that minimizes the MSE for $T=40$ and $r=1$, and the one that minimizes it for $T=\infty$, is small. Similarly, Figure 12 shows that the difference between the constant minimizing the absolute bias for $T=40$ and $r=1$ and the one that minimizes the absolute bias for $T=\infty$ is also small. Thus, both figures indicate that the error introduced by assuming $T=\infty$ when calculating the constants is negligible.
@@ 3.2 Theoretical properties of the estimators
Here we show that, even in the limit as $n\rightarrow\infty$, our estimate achieves better performance than the estimate based on internal branch lengths, $\hat{r}_{Lengths},$ considered in [7].
@@
We will first compute the mean and variance of the reciprocal of the estimator. Recall that $(H_{i}-H_{j})^{+}=\frac{1}{r}(U_{i}-U_{j})^{+}$. Integral calculations in [7] using (2.4) give
@@
$\mbox{Var}\bigg{(}\frac{1}{\hat{r}}\bigg{)}$ $=\mbox{Var}\bigg{(}\frac{1}{(n-1)(n-2)}\sum_{i=1}^{n-1}\sum_{j=1}^{n-1}(H_{i}-H_{j})^{+}\bigg{)}$
@@
The asymptotic normality of $1/\hat{r}$ can be deduced from the fact that $1/\hat{r}$ is a $U$-statistic. Indeed, Theorem 7.1 in [6] states that the distribution of random variables $V$ of the form
@@
from the variance computation above and Theorem 7.1 in [6]. Applying the delta method now yields the conclusion of the lemma. ∎
It remains to show that the estimator $\hat{r}_{Inv}$ has the same asymptotic behavior as $\hat{r}$. We will assume that we have a sample of $n$ individuals taken from a birth and death process at time $T_{n}$. Here $T_{n}$ is finite, so the coalescence times are given by (2.2). We need to assume that $n\rightarrow\infty$, and that $T_{n}\rightarrow\infty$ sufficiently fast relative to $n$. More precisely, we assume, as in [7], that
@@
By Lemmas 4 and 5 in the supplemental material to [7], the random variables can be coupled so that for a sequence of events $(A_{n})_{n=1}^{\infty}$ with $\lim_{n\rightarrow\infty}P(A_{n})=1$, we have
@@
It was shown in [7] that $\sqrt{n}(1/\hat{r}_{Lengths}-1/r)\xRightarrow{n\to\infty}\mathcal{N}(0,1/r^{2}).$ Applying the delta method yields
@@ 3.3 Application to Real Data
We apply $\hat{r}_{MSE},\hat{r}_{Lengths},\hat{r}_{MCMC}$ to data from [16, 15, 3, 10] as saved, e.g. in realCloneData[["cloneTrees"]] of the package cloneRate. These were the same 42 data sets analyzed in [7]. For each of these data sets, the sample size $n$ was between $10$ and $109$. The estimates of the growth rate are shown in Figure 13.
Usually, it holds that $\hat{r}_{MSE}\leq\hat{r}_{MCMC}\leq\hat{r}_{Lengths}$ for the considered data, which is consistent with our simulation results. Exceptions are, for example, PD6629_1 and PD9478_1. In some cases, $\hat{r}_{MCMC}$ and $\hat{r}_{MSE}$ are very similar, e.g. for PD41305 with $n=48.$ In other cases, e.g. AX001 with $n=10$, the differences are larger. As would be expected from the simulation study, the differences between the estimators decreases for increasing sample size $n.$
Figure 13: Application to data from [16, 15, 3, 10]. Red stands for $\hat{r}_{MCMC}$, blue for $\hat{r}_{Lengths}$ and green for $\hat{r}_{MSE}.$ On the $x$-axis are the names of the clone and the sample size and on the $y$-axis are the estimators.
@@ 4 Discussion
We provide an estimator for the growth rate of a birth and death processes which outperforms the state-of-the-art methods ([16, 10, 7]) with respect to the mean squared error and the bias, especially for small sample sizes. We also show analytically that for large sample sizes, our estimator outperforms an estimate based on internal branch lengths from [7].
@@
Our method only applies to exponentially growing populations. However, tumors do not always follow an exponential growth curve ([17]). Nevertheless, as noted in [7], as long as the tumor grows exponentially for a short time, our methods should provide a good estimate of that initial exponential growth rate. Our methods also rely on the assumption that all individuals have the same birth and death rate. This is a limitation because in cell populations, sometimes several driver mutations accumulate, leading to fitness differences among the cells ([13]). However, it was noted by [7] that site frequency spectrum data from the blood cancer subclones studied in [7] agrees well with theoretical predictions for an exponentially growing population in which all cells have the same fitness. Finally, our methods do not take into account any spatial structure in the population and therefore may not be directly applicable to solid tumors.
@@
The code is available on our GitHub website. This code can be used to reproduce the results, i.e. to compare the different estimators for the net growth rate and to apply our method to new data. The code is based on the R-package cloneRate, where the authors also include the real data that has been used in our study.

### v2

@@ 3.1 Simulation Results
We conducted a simulation study to evaluate our estimators $\hat{r}_{MSE}$, $\hat{r}_{Inv}$, and $\hat{r}_{Bias}$, using simulation scenarios similar to those in [Johnson2023]. Specifically, we simulated coalescent point processes for different values of $n$, $r$, and $T$, repeating each scenario 1,000 times for each sample size and growth rate. The simulations were carried out using the R package cloneRate. Following [Johnson2023], we drew $\lambda$ from the uniform distribution on $[r,1+r]$ and subsequently set $\mu=\lambda-r.$ We compared the performance of our estimators to the performance of the estimators $\hat{r}_{MCMC}$, $\hat{r}_{Lengths}$, and $\hat{r}_{MLE}$. Here $\hat{r}_{MCMC}$ is the Markov chain Monte Carlo estimator based on the work of [stadler2009incomplete], which is implemented in cloneRate. Since this is a Bayesian method, its performance will depend on the prior distribution. [williams2022life] proposed uniform priors, and we chose the default prior for $r$ of a uniform distribution on $[0,4]$.
3.1.1 Densities of the estimators
The densities of the six estimators are shown in Figure 3 for $r=0.5$, $T=40$, and $n=5$ and in Figure 3 for $r=0.5$, $T=40$, and $n=20$. These plots show that even for $n=5$, the modes of the densities of $\hat{r}_{Lengths}$, $\hat{r}_{MLE}$, $\hat{r}_{MCMC}$, and $\hat{r}_{Inv}$ are similar to the true growth rate. According to the density plots, $\hat{r}_{Bias}$ and especially $\hat{r}_{MSE}$ typically underestimate the growth rate for small $n$. Moreover, the densities are all very skewed for $n=5$.
For $n=20,$ the differences among the estimators are much smaller than for $n=5.$ The skewness decreases with increasing sample size so that for $n=20$, the densities look similar to the density of a normal distribution. This aligns with results in [Johnson2023].
To understand Figure 3, observe that if all of the coalescence times $H_{i,n}$ coincidentally happen to be very close to one another, the denominator in (2.8) will be small, making the estimate $\hat{r}$ very large. Although this is exceptionally rare when $n=20$, and quite rare even when $n=5$, it happens often enough when $n=5$ to have a substantial influence on the mean squared error, and to a lesser extent the bias, of the estimate. Therefore, a small value of $c(n)$, which mitigates the effect of these rare events, will minimize the mean squared error. Consequently, the estimator $\hat{r}_{MSE}$ that minimizes the mean squared error often underestimates $r$ by nearly a factor of $2$ when $n$ is small.
3.1.2 Mean Squared Error, Mean Absolute Error, and Bias
We compared the root mean squared error (RMSE), the mean absolute error, and the bias for the six different estimators when $r=0.5$ and $T=40$, and when $r=1$ and $T=40$, for several different values of $n$. The root mean squared errors are shown in Figures 5 and 5. These figures show that the RMSE of the estimator $\hat{r}_{{MSE}}$ is the lowest across all sample sizes. The differences among the estimators decrease as $n$ increases.
Moreover, not only does $\hat{r}_{{MSE}}$ outperform $\hat{r}_{{MCMC}},\hat{r}_{{MLE}}$ and $\hat{r}_{{Lengths}}$ on this criterion, but also $\hat{r}_{{Bias}}$ and $\hat{r}_{{Inv}}$ achieve lower RMSEs than these previously existing methods when $n\geq 7$ in the case $r=0.5$ and when $n\geq 9$ in the case $r=1$.
Figure 4: Comparison of the RMSE of the six estimators when $r=0.5$ and $T=40$.
Figure 5: Comparison of the RMSE of the six estimators when $r=1$ and $T=40$.
@@ 3.1.2 Mean Squared Error, Mean Absolute Error, and Bias
The mean absolute errors of the estimators are shown in Figures 7 and 7. Here, $\hat{r}_{Bias}$ has the lowest mean absolute error across the six estimators. However, $\hat{r}_{MSE}$ and $\hat{r}_{Inv}$ also perform better than $\hat{r}_{MCMC},\hat{r}_{MSE}$ and $\hat{r}_{Lengths}$ for most values of $n$, $r$, and $T$.
@@
The bias of the estimators is presented in Figures 9 and 9. According to the figures, the estimates $\hat{r}_{Lengths}$, $\hat{r}_{MLE}$, and $\hat{r}_{MCMC}$ all tend to overestimate the growth rate on average for small $n$. The estimate $\hat{r}_{Bias}$, which was designed to be exactly unbiased when $T=\infty$, remains nearly unbiased even when $T=40$. Note that the estimate is nearly unbiased even though the mode of the density is considerably below the true value of $r$ because occasional large overestimates of $r$ affect the mean. The estimate $\hat{r}_{Inv}$ has a slight upward bias, as would be expected from Jensen’s inequality because it is the reciprocal of an estimator which is unbiased for $1/r$ when $T=\infty$. The estimate $\hat{r}_{MSE}$ underestimates the growth rate on average when $n$ is small. It acts as a shrinkage estimator, minimizing the MSE by introducing a downward bias to reduce the variance and mitigate the impact of rare events in which the growth rate is vastly overestimated.
@@
We also compared the results when $n=5$ and $T=100$, and when $n=20$ and $T=100$, for several different values of $r$. These results are shown in Figures 11, 11, 13, and 13. These figures illustrate that the ratios $\mathrm{RMSE}/r$ and $\mathrm{MAE}/r$ for $\hat{r}_{{MSE}}$ and $\hat{r}_{{Inv}}$ remain roughly constant across the different values of $r$. On the other hand, the ratios $\mathrm{RMSE}/r$ and $\mathrm{MAE}/r$ for $\hat{r}_{{MCMC}}$ tend to decrease as $r$ increases. Consequently, by these criteria, the estimators $\hat{r}_{{MSE}}$ and $\hat{r}_{{Inv}}$ perform better for small $r$, while $\hat{r}_{{MCMC}}$ performs better for large $r$.
This behavior can be explained by the fact that the prior distribution imposes an upper bound of 4 on $\hat{r}_{{MCMC}}$, as this is the default value in the software package. As a result, the performance of $\hat{r}_{{MCMC}}$ is less impacted by rare events in which the other methods estimate extremely high values of $r$. The impact of this is discussed in more detail in Section 5.5 in the appendix. It is shown in Figure 20 in the appendix that the choice of the maximal growth rate substantially impacts the mean squared error of the estimator $\hat{r}_{MCMC}$. We believe that one of the benefits of our methods is that they do not require specifying a prior distribution in advance.
Figure 10: Comparison of the RMSE/r of three estimators when $n=5$ and $T=100$.
Figure 11: Comparison of the RMSE/r of three estimators when $n=20$ and $T=100$.
Figure 12: Comparison of the MAE/r of three estimators when $n=5$ and $T=100$.
Figure 13: Comparison of the MAE/r of three estimators when $n=20$ and $T=100$.
We also compared different values of $T$. For example, when comparing Figures 11 and 11 with Figure 5, we observe relatively little difference in performance between $T=100$ and $T=40$. Overall, the estimators $\hat{r}_{MSE}$, $\hat{r}_{Inv}$, and $\hat{r}_{Bias}$ achieve good performance relative to other estimators. They have the computational advantage that they do not require Markov Chain Monte Carlo, and they do not require the specification of a prior. Which of the estimators is best will depend on the performance criteria. The estimator $\hat{r}_{MSE}$ achieves the lowest mean squared error, while $\hat{r}_{Bias}$ has the advantage of being nearly unbiased. On the other hand, both of these estimators tend to underestimate $r$ most of the time in small samples. The estimator $\hat{r}_{Inv}$ appears to be a good compromise. It still achieves a lower mean absolute error than $\hat{r}_{MCMC}$ in most simulations, and the mode of the distribution is close to the true growth rate. It also has the benefit that $c_{Inv}(n)$ can be computed exactly.
@@ 3.1.4 Influence of the Approximation $T=\infty$ on the Quality of the Estimator
Now, we quantify the error introduced by this approximation. Given $n$, $T$, and $r$, we can determine by simulation, for any value of the constant $c$, the root mean squared error and the absolute bias of the resulting estimator of $r$. In Figures 16 and 16, we plot the RMSE and the bias as a function of the constant.
Figure 15: Influence of the constant on the RMSE for $T=40$, $n=10$, and $r=0.5$.
Figure 16: Influence of the constant on the MAE for $T=40$, $n=10$, and $r=0.5$.
The red dotted line in Figure 16 lies close to the minimum point of the curve, which means that the difference between the constant that minimizes the MSE for $T=40$ and $r=1$, and the one that minimizes it for $T=\infty$, is small. Similarly, Figure 16 shows that the difference between the constant minimizing the absolute bias for $T=40$ and $r=1$ and the one that minimizes the absolute bias for $T=\infty$ is also small. Thus, both figures indicate that the error introduced by assuming $T=\infty$ when calculating the constants is negligible.
@@ 3.2 Theoretical properties of the estimators
Here we show that, even in the limit as $n\rightarrow\infty$, our estimate achieves better performance than the estimate based on internal branch lengths, $\hat{r}_{Lengths},$ considered in [Johnson2023].
@@
We will first compute the mean and variance of the reciprocal of the estimator. Recall that $(H_{i}-H_{j})^{+}=\frac{1}{r}(U_{i}-U_{j})^{+}$. Integral calculations in [Johnson2023] using (2.4) give
@@
[…]
@@
The asymptotic normality of $1/\hat{r}$ can be deduced from the fact that $1/\hat{r}$ is a $U$-statistic. Indeed, Theorem 7.1 in [hoeffding1992class] states that the distribution of random variables $V$ of the form
@@
from the variance computation above and Theorem 7.1 in [hoeffding1992class]. Applying the delta method now yields the conclusion of the lemma. ∎
It remains to show that the estimator $\hat{r}_{Inv}$ has the same asymptotic behavior as $\hat{r}$. We will assume that we have a sample of $n$ individuals taken from a birth and death process at time $T_{n}$. Here $T_{n}$ is finite, so the coalescence times are given by (2.2). We need to assume that $n\rightarrow\infty$, and that $T_{n}\rightarrow\infty$ sufficiently fast relative to $n$. More precisely, we assume, as in [Johnson2023], that
@@
By Lemmas 4 and 5 in the supplemental material to [Johnson2023], the random variables can be coupled so that for a sequence of events $(A_{n})_{n=1}^{\infty}$ with $\lim_{n\rightarrow\infty}\mathbb{P}(A_{n})=1$, we have
@@
It was shown in [Johnson2023] that $\sqrt{n}(1/\hat{r}_{Lengths}-1/r)\xRightarrow{n\to\infty}\mathcal{N}(0,1/r^{2}).$ Applying the delta method yields
@@ 3.3 Application to Real Data
We apply $\hat{r}_{MSE}$, $\hat{r}_{Inv}$, and $\hat{r}_{MCMC}$ to data from [williams2022life, van2021reconstructing, fabre2022longitudinal, mitchell2022clonal] as saved, e.g., in realCloneData[["cloneTrees"]] of the package cloneRate. These were the same 42 data sets analyzed in [Johnson2023]. For each of these data sets, the sample size $n$ was between $10$ and $109$. The estimates of the growth rate are shown in Figure 17.
Usually, it holds that $\hat{r}_{MSE}\leq\hat{r}_{Inv}\leq\hat{r}_{MCMC}$ for the considered data. For most of the data sets, the three estimators give very similar results, although there are a few exceptions where there are larger differences among the methods. As would be expected from the simulation study, the differences among the estimators decreases for increasing sample size $n.$ Overall, we see that, much like the methods proposed by [Johnson2023], the methods $\hat{r}_{Inv}$ and $\hat{r}_{MSE}$ give results on these real data sets that are comparable to those obtained from Markov chain Monte Carlo methods, for less computational cost.
Figure 17: Application to data from [williams2022life, van2021reconstructing, fabre2022longitudinal, mitchell2022clonal]. On the $x$-axis are the names of the clone and the sample size and on the $y$-axis are the estimators.
@@ 4 Discussion
We provide an estimator for the growth rate of a birth and death processes which outperforms the state-of-the-art methods ([williams2022life, mitchell2022clonal, Johnson2023]) with respect to the mean squared error and the bias, especially for small sample sizes. We also show analytically that for large sample sizes, our estimator outperforms an estimate based on internal branch lengths from [Johnson2023].
@@
Our method only applies to exponentially growing populations. However, tumors do not always follow an exponential growth curve ([zahid2021forecasting]). Nevertheless, as noted in [Johnson2023], as long as the population grows exponentially for a short time, our methods should provide a good estimate of that initial exponential growth rate because the coalescence times of the sampled lineages are usually early in the history of the population. Mathematical results in this direction were established by [cheek2022coalescent]. To test the assumption of an exponentially growing population, one could use skyline methods as introduced by [pybus2000integrated], [drummond2005], and [stadlerskyline]. Skyline methods can reconstruct the population size history under different demographic models.
Our methods also rely on the assumption that all individuals have the same birth and death rates. This is a limitation because in cell populations, sometimes several driver mutations accumulate, leading to fitness differences among the cells ([scott2017somatic]). However, it was noted by [Johnson2023] that site frequency spectrum data from the blood cancer subclones studied in [Johnson2023] agrees well with theoretical predictions for an exponentially growing population in which all cells have the same fitness. Finally, our methods do not take into account any spatial structure in the population and therefore may not be directly applicable to solid tumors.
@@
The code is available on our GitHub website (https://github.com/CarolaHeinzel/Estimation_NetGrowthRate). This code can be used to reproduce the results, i.e. to compare the different estimators for the growth rate and to apply our method to new data. The code is based on the R-package cloneRate, where the authors also include the real data that has been used in our study.